Liquidating Partnerships

1. Prior to liquidating their partnership, MacPherson and Gentry had capital accounts of $45,000 and $76,000, respectively. Prior to liquidation, the partnership had no cash assets other than what was realized from the sale of assets. These partnership assets were sold for $116,000. The partnership had $5,000 of liabilities. MacPherson and Gentry share income and losses equally.

Determine the amount received by MacPherson as a final distribution from liquidation of the partnership.
$

2.

Liquidating Partnerships—Deficiency

Prior to liquidating their partnership, Jolly and Russo had capital accounts of $29,000 and $108,000, respectively. The partnership assets were sold for $51,000. The partnership had no liabilities. Jolly and Russo share income and losses equally.

Required:

a. Determine the amount of Jolly's deficiency.
$

b. Determine the amount distributed to Russo, assuming Jolly is unable to satisfy the deficiency.
$
Answer #1

1.

Asset Prior to liquidation (45,000+76,000+5,000) 126,000
Partnership asset sold 116,000
Profit on sales 10,000
Gentry Share of profit 5,000
Amount Received by Gentry:
Capital Balance 76,000
Gentry Share of profit on liquidation 5,000
81,000

2.

a. Amount of Jolly Deficiency

Total Capital = Jolly Capital+Russo

= 29,000 + 108,000

= 137,000

Total assets = 51,000

Loss on Dissolution = Total Capital - Total assets

= 137,000-51000

= 86,000

Jolly part of loss = 43,000(86,000/2)

Amount of Jolly Deficiency = Jolly part of Loss - Jolly Capital balance

= 43,000-29,000

= 14,000

b. Amount distributed to russo is whole $ 51,000 realized from the assets because Jolly capital gone negative due to Loss share and he is not eligible for the single dollar payment.